OpenStudy (anonymous):

A circle has an area of 324╥ cm². What is the radius? A) 18 c, B) 36╥ cm C) 18╥ cm D) 36 cm (I think It's D) I got d=²r ... d = (2)18 .. BUT I'M NO SURE IF THAT

Parth (parthkohli):

Then: \(\Large \color{Black}{\Rightarrow \pi r^2 = 324 }\) Solve.

Parth (parthkohli):

Actually: \(\Large \color{Black}{\Rightarrow \pi r^2 = 324\pi}\)

OpenStudy (anonymous):

I got 18 x 18 = 324

OpenStudy (anonymous):

area = (pi) r^2 324/(pi) = r^2 \[\sqrt{\frac{324}{\pi}} = r\]

OpenStudy (anonymous):

so then I did 2(18) = 36

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow r^2 = 324 }\) \(\Large \color{Black}{\Rightarrow r = \sqrt{324} = 18 }\) Diameter is 2(18) = 36

Parth (parthkohli):

And the radius is 18

OpenStudy (anonymous):

So the diameter is 36? Radius is 18?

OpenStudy (anonymous):

ooooh, area is 324xpi though I was going mad. That's a strange pi in the question !

OpenStudy (anonymous):

aahh I only help ALT + 210 to get that pi

OpenStudy (anonymous):

the aqn editor has also gone down hill here used to work ok!